UT Jun 7 2022 - DEC Night



Night Info

  • Sunset 19:25
  • Nautical Twilight Ends 20:28
  • Astro Twilight Ends 21:06
  • Night Center 00:18
  • Moon sets 00:56
  • Astro Twilight Starts 03:31
  • Nautical Twilight Starts 04:08
  • Sunrise 05:11

Staff

  • OSA: S.Allanson
  • ISp: Mtn Staff
  • Mtn Manager: R.Hansen
  • ISA/Observer: B.Rothberg/J.Power
  • AO: J.Power

Night Plan

  • DONE Start LUCI-1 WFS in afternoon, install software update verify works in SL
  • DONE TCS/AOS patch
  • DONE DMS test of increasing telemetry from 2Hz to 5Hz (Bec)
    • [JMH: This is verification that all the rest of TCS is still working fine, so does not need explicit time on-sky.]
    • Completed early in day
  • Starting Configuration bino-MODS
  • Science Time
    • DONE SpecPhot Std - Feige 34 in twilight (as soon as possible) - Dual Grating (script in Rothberg OT file)
    • DONE mid-z ULIRG target - Rothberg - 2hr - target hits limit at 23:15hr
      • Seeing < 1.2"
      • thin clouds OK
      • FF1106 - blind offset - twilight-23:06
      • OT: /home/lbto/Service_Observing/2022A_D_E/MODS/Rothberg/midzUlirgs/midz_ULIRGs_V17.xml
    • DONE Odd Radio Circle target - Becker - 2.5hr - target hits limit at 02:45hr
      • Clear
      • Seeing < 1.2"
      • Grey skies
      • RA: 14:21 DEC 35:14 - blind offset - 23:07-01:37hr
      • OT: /home/lbto/Service_Observing/2022A_D_E/MODS/Becker/ORC_Hunter.xml
  • DONE Reconfigure for bino LUCI/LUCI-1 authorized - 01:40-01:50hr
  • AO engineering tests
    • OT File: /home/lbto/Service_Observing/2022A_D_E/LUCI/AO_ESM/AO_Testing_V89.xml
    • DONE -ish - OSCO - check Bin 1 stability 01:50-02:10hr
    • DONE setz0 test
      • Verify first half of night that no impact to SL observations
      • verify NCPA offsets are applied visually with get_z0
        • verification with GOPT GUI
      • verify no BCU communication issues
      • Bin 2 target (R~12-12.5 - WFSmag~13-13.5) - with dithers and off-field dithers - 02:10hr-03:10hr
        • Target TBD (backup target TBD)
        • run with NCPA off as a sanity check/baseline - 0.5hr
        • run with NCPA on - 0.5hr
        • Visible Bin 2 targets
          • RA 16-17hr Folder, Star 4 WD 1608+477 R=11.8 (predicted WFSmag = 13) - visible all night
          • RA 16-17hr Folder, Star 5 SBSS 1709+535 R=12.8 (predicted WFSmag 14) - visible all night
          • RA 16-17hr Folder, Star 6 PG 1654+138 R=13 (predicted WFSmag = 14.2) - visible all night
          • RA 18-19hr Folder Star 2 HD 349137 (Dec +18) R=11.7 (predicted WFSmag = 12.9) - visible 21:30-twilight
          • RA 18-19hr Folder Star 6 HD 333105 (Dec +30) R=11.8 (predicted WFSmag = 13.0) - visible 22:30-twilight
      • DONE Also Bin 3 and Bin 1
    • ESM characterization - 03:01hr-twilight - or other tasks if higher priority
      • Dual BLR - mixed mode ESM and SL
      • Scripts: /home/lbto/Service_Observing/2022A_D_E/LUCI/DualBLR/SDSSJ1713+3256_ESM_SL.xml
        • seeing up to 1.2" thin clouds
        • 0.75hr

Poor Weather Backup

  • Bino-LUCIs
    • Local Mergers - Rothberg
      • Conditions: If you can guide you can observe:
      • UGC 9829 - RA 15hr Dec -01 - visible twilight-02hr
      • 1.5hr-2hr depending on conditions
      • OT: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V50.xml
    • ICRF3 project
      • OT: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/ICRF3/ICRF3_LUCI_V25_LBTO.xml
      • Pick based on visibility and priority
    • Photometric Standards
      • Must be clear - any seeing
      • If seeing so terrible can't guide, then export as track mode
      • OT: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards/LUCI_Phot_ZP_OT_V4.xml
    • ESM testing in poor conditions
      • Can be run to characterize performance in poor conditions (contined testing)
      • Runs through K,H,J - dithers in and out of FOV
      • Parent Directory: /home/lbto/Service_Observing/2022A_D_E/LUCI/AO_ESM/ESM
        • Mixed mode scripts have already been generated and updated to do mixed mode
        • M92 - bright (R=9.4) and faint (R=15.2)- 1hr each - visible all night
        • M52 - R=10.7 = visible 01hr-twilight
  • Bino-MODS
    • Local Mergers - Rothberg
      • Conditions - Up to 2" seeing, 1-2 mags extinction
      • IRAS 21219-1757 - 2.5hr w/overheads
      • OT: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V50.xml
      • Scripts made for any combo of missing MODS channels
      • visible 02:30hr-twlight - observe SpecStd before - get as many photons as can before twilight too bright
    • Spectrophotometric Standard in all configuations
      • Must be clear, any seeing

Night Log

Nightlog located in Queue and Service Observing -> ECD Nights -> ECD 2022 Shared Drive

https://docs.google.com/document/d/1hO-Lw8ZYL1oIWYht1xh_uV0NgCNVizxN6TC2M3Gfols/edit?usp=sharing

Closed Dome

-- BarryRothberg - 03 Jun 2022
Topic revision: r10 - 07 Jun 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback